# ForceMimic: Force-Centric Imitation Learning with Force-Motion Capture System for Contact-Rich Manipulation

### v2

@@ Abstract
In most contact-rich manipulation tasks, humans apply time-varying forces to the target object, compensating for inaccuracies in the vision-guided hand trajectory. However, current robot learning algorithms primarily focus on trajectory-based policy, with limited attention given to learning force-related skills. To address this limitation, we introduce ForceMimic, a force-centric robot learning system, providing a natural, force-aware and robot-free robotic demonstration collection system, along with a hybrid force-motion imitation learning algorithm for robust contact-rich manipulation. Using the proposed ForceCapture system, an operator can peel a zucchini in 5 minutes, while force-feedback teleoperation takes over 13 minutes and struggles with task completion. With the collected data, we propose HybridIL to train a force-centric imitation learning model, equipped with hybrid force-position control primitive to fit the predicted wrench-position parameters during robot execution. Experiments demonstrate that our approach enables the model to learn a more robust policy under the contact-rich task of vegetable peeling, increasing the success rates by 54.5% relatively compared to state-of-the-art pure-vision-based imitation learning. Hardware, code, data and more results would be open-sourced on the project website at https://forcemimic.github.io.
@@ I Introduction
Humans can take use of force sensing, fine muscle force control to achieve better manipulations, from grasping [1], lifting [2] to peeling [3]. The exploitation of force can detect and correct the error brought by vision-based motion planning.
Inspired by these neuroscience results, we want to explore the utility of force in robot learning. However, the force-centric manipulation demonstration data is hard to collect. Substantial human videos exist on the Internet, but no interaction force data is recorded. Teleoperation [4] is a popular data collection approach, enabling operators remotely to control robot finishing manipulation tasks. Particularly, force-feedback teleoperation shows a potential path to the force-centric data collection. But it cannot give the operator a natural manipulation experience, harmful to smooth action execution and precise force control. Recently, portable handheld devices [5, 6] make in-the-wild learning possible. They make use of SLAM tracking camera to record human hand or handheld gripper pose trajectory. In addition to the removal of real robot, it provides an additional advantage as almost direct interaction between human and objects, which is critical in contact-rich force-centric manipulation.
@@
To handle the above challenges, we propose ForceMimic, a force-centric robot learning system, providing natural, force-aware and robot-free robotic demonstration collection experience and force-centric imitation learning algorithm equipped with hybrid force-position control to achieve robust contact-rich manipulation (see Fig. LABEL:fig:teaser). We first develop ForceCapture, a handheld robot-free data collection system to record high-quality pure interaction wrench, by ratchet locking and gravity compensation. After that, HybridIL leverages the data to train a force-aware policy that outputs wrench-position parameters. And we apply hybrid force-position control to fit not only predicted pose trajectory but also the predicted force parameters, achieving robust execution against error-prone visual guide. Using ForceCapture, operators can peel a zucchini in just 5 minutes, while force-feedback teleoperation takes over 13 minutes and struggles with smooth peeling. Robot experiments show that HybridIL achieves a 54.5% higher success rate during contact-rich peeling compared to state-of-the-art vision-based imitation learning.
@@ II Related Work
Force perception and control plays a crucial role in manipulation tasks, providing valuable and complementary information with visual guidance [30]. Several works have explored the force leverage in contact-rich robotic manipulation, ranging from opening bottle caps [31], assembling [32] to playing Jenga [33]. Recently, MOMA-Force [34] utilizes the visual representation similarity to retrieve target action and wrench from the expert database and uses a PID-based controller [35, 36] to control the robot. In this paper, We propose a new paradigm of using orthogonal hybrid force-position control primitives to fit the model’s predicted wrench-position parameters.
@@ III Method
ForceMimic first employs ForceCapture, a handheld robot-free data collection system (detailed in Sec. III-A), to naturally gather force-centric human demonstration data. Then, we transfer the robot-free data to (pseudo-)robot data (detailed in Sec. III-B), migrating the domain gap. Leveraging this data, HybridIL learns to predict wrench-pose trajectory, and applies hybrid force-position control to fit the predicted force-position parameters (detailed in Sec. III-C), enabling robust performance in contact-rich manipulation tasks. The overall pipeline is illustrated in Fig. 1.
Figure 1: Overview of the pipeline. (a) We first transfer the collected robot-free data to (pseudo-)robot data, migrating the domain gap. The captured wrench is compensated to account for self-gravity effects. The pose recorded by SLAM camera is transformed as the robot TCP pose. And RGB-D observation images are backprojected into point cloud and filtered out unrelated points. (b) Leveraging this data, a diffusion-based policy is learned, with both pose and wrench predicted, conditioned on the encoded point cloud features, history pose and diffusion timestep embeddings. (c) According to the predicted force value, either IK joint position primitive or hybrid force-position primitive is selected, and fits the output force-position parameters to conduct execution actions.
@@ III-A Hardware Design: ForceCapture
Accurately, naturally, and cost-effectively capturing force data during contact-rich manipulation remains a significant challenge. Different data collection methods can lead to varying force interaction dynamics, where even identical trajectories may involve different interaction forces. Therefore, preserving the natural force interaction during object manipulation is crucial for designing an effective force-position data collection system. Inspired by existing handheld motion data collection devices [5, 6], we developed a low-cost, versatile, and robot-free force-position capture device, ForceCapture. In designing this device, we consistently adhered to the following objectives:
@@
The overall design is shown in Fig. 2, which illustrates two versions, one with a fixed tool and the other with an adaptive gripper. At its core, both designs share the feature of a six-axis force sensor placed between the end-effector and the user’s gripping handle, which can be used to capture the effector-environment interaction forces. Additionally, a SLAM camera positioned near the center of the force sensor records the motion data during interaction. The user grips the handle to directly operate the tool or control the fingers for grasping and manipulation tasks. The rack-and-pinion mechanism of the gripper version at the base of both fingers ensures synchronized movement of the grippers. The pinion is connected to an encoder, which records the opening distance of the grippers. The continuous width value is determined based on the calibrated relationship between the encoder angle and the gripper’s width.
@@ III-B Data Collection and Transfer
ForceCapture is designed to record only the interaction forces between the end-effector and the external environment. However, the force sensor measures the combined forces, including the tool’s gravitational and inertial forces. Therefore, it is necessary to subtract the external forces generated by the tool or gripper from the force sensor data. We assume that the data collection process with ForceCapture is quasi-static, meaning that at each position, the forces are in static equilibrium, and we only need to compensate for the tool’s gravity. To perform the gravity compensation, we first move ForceCapture in a quasi-static manner for a certain period while recording the pose and wrench data. Using the static equilibrium forces at each position, we construct an overdetermined system of equations to estimate the tool’s center of mass and weight using least-squares solution.
@@ III-C Learning Algorithm: HybridIL
It is important to note that wrench and position control must be orthogonal. While our model does not explicitly model the orthogonality of wrench and position, we achieve this through an orthogonal force-position hybrid controller that aligns with the model’s predicted force-position parameters. This approach differs from conventional imitation learning methods, which typically use a fixed lower-level position controller to track the position commands prediction by the model. HybridIL employs two distinct control primitives to fit the model’s predicted force-position parameters, demonstrated in Fig. 1 (c). When the predicted force is below a threshold of 6N, an IK-based joint position controller is used. If the predicted force exceeds 6N over consecutive steps, a hybrid force-position controller is employed to execute the model’s predicted parameters. The orthogonal force-position matching approach is illustrated in Fig. 3. For force-position actions where the force exceeds 6N continuously, the motion direction is determined based on the positional information before and after. The corresponding predicted force information is projected onto the orthogonal plane of the motion direction, which defines the force control parameters during execution. For the initial step of hybrid force-position control, if the end-effector has not yet made contact with the object, a pressing control in the opposite direction of force control is applied to achieve stable contact. These functionalities are realized using Flexiv RDK of joint position control and hybrid force-position control primitives to execute the force-position actions of HybridIL.
@@ IV Experiments
Figure 4: Experimental setup for data collection efficiency comparison and the time required to fully peel a zucchini.
@@ Results
The results of the four methods are summarized in TABLE I, and the peeled skins are shown in Fig. 5. The Raw DP method achieved a motion success rate of 80% (16/20), with instances of failure detailed in the middle of Fig. 5. Failures marked as ② involved excessive force during the peeling process, which resulted in damage to the zucchini. One instance even broke the bottom of the zucchini, as shown in the bottom middle of Fig. 5. Failure marked as ④ resulted from no contact with the zucchini, hence no peeling occurred. In contrast, HybridIL demonstrated a 100% success rate (20/20), with all attempts resulting in successful contact and peeling, as illustrated in the left of Fig. 5. When the success criterion was increased to a continuous peeling length of more than 10 cm, both Raw DP and HybridIL experienced a decrease in success rates. Raw DP’s success rate dropped to 55%, with additional failure cases marked as ① and ③. Case ① indicated peeling lengths shorter than 10 cm, and case ③ involved peeling breakage, attributed to discontinuity between the output poses, which caused peeling interruptions. For HybridIL, the success rate decreased to 85%, with failures in cases ① and ③. These failures were due to the premature ending of the output force-position parameters, leading to an early switch from the hybrid force control primitives to IK-based joint position control primitives, which caused peeling discontinuities.
@@
Figure 5: Visualization of the peeled skins.
Figure 6: Force curves during different manipulation.
The models that incorporate force as an input, including Force DP and Force+Hybrid DP, performed poorly. While the initial motion leading to the contact with the zucchini was generally correct, once contact occurred, these models struggled to predict the correct pose and force, making successful peeling nearly impossible. This result is counterintuitive—–one would expect that adding force sensing would improve peeling performance, but instead, it worsened the outcome. The reasons for this can be understood from the interactive force curves of the Raw DP and HybridIL during the peeling process, as shown in Fig. 6. Although Raw DP successfully peeled the zucchini, the interactive forces were significantly higher, averaging around 20N and reaching over 40N in some areas. In contrast, the dataset from which these models were trained exhibited much lower interaction forces, around 6N as shown in the left of Fig. 6. This mismatch between the input forces and the force distribution in the dataset made it difficult for the models to predict the correct actions. The inconsistency between the force interaction controller used during robot deployment and data collection might be a potential factor contributing to the issue. Addressing this discrepancy could improve the model’s performance. Effectively utilizing the force data collected by ForceCapture as sensory input remains an open challenge and a promising direction for future research. Further exploration on how to better integrate force perception into control strategies could lead to significant advancements in improving task performance. Additionally, as shown in Fig. 5 and Fig. 6, the HybridIL method maintained an average interaction force of 7.5N, closely aligning with the model’s predicted forces, resulting in evenly peeled sections with consistent thickness and width. However, a slight deviation from the force distribution in the dataset was still noticeable, which likely explains why the Force+Hybrid DP failed to yield good results. Nonetheless, Force+Hybrid DP demonstrated an improvement over the Force DP.
The above experiments demonstrate that using hybird force-position control primitives to fit the predicted force-position parameters is effective, further validating the ForceMimic system. Additional experimental results can be found on the project website.
@@ V Conclusion and Discussions
We provide a force-position capture system for contact-rich manipulations and offer an initial exploration of learning human force-position skills. However, there is still room for improvement. 1) Our model uses a simple MLP to represent point clouds, robot pose, and force. In the future, we could explore more advanced multimodal representations that combine visual, force, and robot state data to improve the model’s generalization to diverse skills. 2) HybridIL only employs two control primitives to fit the model’s predicted force-position parameters. Future research could involve exploring more control primitives to better align with model outputs, and the model itself could potentially predict the most suitable primitive and corresponding parameters in advance. 3) ForceMimic has so far demonstrated success with a single peeling skill. In the future, the system could be extended to more force-centric tasks such as scraping, wiping, or inserting, as well as combining multiple skills for long-horizon tasks. For example, the complete peeling of a zucchini could include steps like peeling, grasping, and flipping.

### v3

@@ Abstract
In most contact-rich manipulation tasks, humans apply time-varying forces to the target object, compensating for inaccuracies in the vision-guided hand trajectory. However, current robot learning algorithms primarily focus on trajectory-based policy, with limited attention given to learning force-related skills. To address this limitation, we introduce ForceMimic, a force-centric robot learning system, providing a natural, force-aware and robot-free robotic demonstration collection system, along with a hybrid force-motion imitation learning algorithm for robust contact-rich manipulation. Using the proposed ForceCapture system, an operator can peel a zucchini in 5 minutes, while force-feedback teleoperation takes over 13 minutes and struggles with task completion. With the collected data, we propose HybridIL to train a force-centric imitation learning model, equipped with hybrid force-position control primitive to fit the predicted wrench-position parameters during robot execution. Experiments demonstrate that our approach enables the model to learn a more robust policy under the contact-rich task of vegetable peeling, increasing the success rates by 54.5% relatively compared to state-of-the-art pure-vision-based imitation learning. Hardware, code, data and more results can be found on the project website at https://forcemimic.github.io.
@@ I Introduction
Humans can take use of force sensing, fine muscle force control to achieve better manipulations, from grasping [1], lifting [2] to peeling [3]. The exploitation of force can detect and correct the error brought by vision-based motion planning. Inspired by these neuroscience results, we want to explore the utility of force in robot learning. However, the force-centric manipulation demonstration data is hard to collect. Substantial human videos exist on the Internet, but no interaction force data is recorded. Teleoperation [4] is a popular data collection approach, enabling operators remotely to control robot finishing manipulation tasks. Particularly, force-feedback teleoperation shows a potential path to the force-centric data collection. But it cannot give the operator a natural manipulation experience, harmful to smooth action execution and precise force control. Recently, portable handheld devices [5, 6] make in-the-wild learning possible. They make use of SLAM tracking camera to record human hand or handheld gripper pose trajectory. In addition to the removal of real robot, it provides an additional advantage as almost direct interaction between human and objects, which is critical in contact-rich force-centric manipulation.
@@
To handle the above challenges, we propose ForceMimic, a force-centric robot learning system, providing natural, force-aware and robot-free robotic demonstration collection experience and force-centric imitation learning algorithm equipped with hybrid force-position control to achieve robust contact-rich manipulation (see Fig. LABEL:fig:teaser). We first develop ForceCapture, a handheld robot-free data collection system to record high-quality pure interaction wrench, i.e. combination of force and torque, by ratchet locking and gravity compensation. After that, HybridIL leverages the data to train a force-aware policy that outputs wrench-position parameters. And we apply hybrid force-position control to fit not only predicted pose trajectory but also the predicted force parameters, achieving robust execution against error-prone visual guide. Using ForceCapture, operators can peel a zucchini in just 5 minutes, while force-feedback teleoperation takes over 13 minutes and struggles with smooth peeling. Robot experiments show that HybridIL achieves a 54.5% higher success rate during contact-rich peeling compared to state-of-the-art vision-based imitation learning.
@@ II Related Work
Force perception and control plays a crucial role in manipulation tasks, providing valuable and complementary information with visual guidance [30]. Several works have explored the force in contact-rich robotic manipulation, ranging from opening bottle caps [31], assembling [32] to playing Jenga [33]. Recently, MOMA-Force [34] utilizes the visual representation similarity to retrieve target action and wrench from the expert database and uses a PID-based controller [35, 36] to control the robot. ForceSight [37] presents a transformer-based robotic planner that generates force-based objectives given a text input and an RGBD image. In this paper, We propose a new paradigm of using orthogonal hybrid force-position control primitives to fit the model’s predicted continue wrench-position parameters.
@@ III Method
ForceMimic first employs ForceCapture, a handheld robot-free data collection system (detailed in Sec. III-A), to naturally gather force-centric human demonstration data. Then, we transfer the robot-free data to (pseudo-)robot data (detailed in Sec. III-B), bridging the domain gap. Leveraging this data, HybridIL learns to predict wrench-pose trajectory, and applies hybrid force-position control to fit the predicted force-position parameters (detailed in Sec. III-C), enabling robust performance in contact-rich manipulation tasks. The overall pipeline is illustrated in Fig. 1.
Figure 1: Overview of the pipeline. (a) We first transfer the collected robot-free data to (pseudo-)robot data, bridging the domain gap. The captured wrench is compensated to account for self-gravity effects. The pose recorded by SLAM camera is transformed as the robot TCP pose. And RGB-D observation images are backprojected into point cloud and filtered out unrelated points. (b) Leveraging this data, a diffusion-based policy is learned, with both pose and wrench predicted, conditioned on the encoded point cloud features, history pose and diffusion timestep embeddings. (c) According to the predicted force value, either IK joint position primitive or hybrid force-position primitive is selected, and fits the output force-position parameters to conduct execution actions.
@@ III-A Hardware Design: ForceCapture
Accurately, naturally, and cost-effectively capturing force data during contact-rich manipulation remains a significant challenge. Inspired by existing handheld motion data collection devices [5, 6], we developed a low-cost, versatile, and robot-free force-position capture device, ForceCapture. To design ForceCapture, we consistently adhered to the following objectives:
@@
The overall design is shown in Fig. 2, which illustrates two versions, one with a fixed tool and the other with an adaptive gripper. At its core, both designs share the feature of a six-axis force sensor placed between the end-effector and the user’s gripping handle, which can be used to capture the effector-environment interaction wrench. Additionally, a SLAM camera positioned near the center of the force sensor records the motion data during interaction. The user grips the handle to directly operate the tool or control the fingers for grasping and manipulation tasks. The rack-and-pinion mechanism of the gripper version at the base of both fingers ensures synchronized movement of the grippers. The pinion is connected to an encoder, which records the opening distance of the grippers. The continuous width value is determined based on the calibrated relationship between the encoder angle and the gripper’s width.
@@ III-B Data Collection and Transfer
ForceCapture is designed to record only interaction forces between end-effector and external environment. However, the force sensor measures the combined forces, including the tool’s gravitational and inertial forces. Therefore, it is necessary to subtract the external forces generated by the tool or gripper from the force sensor data. We assume that the data collection process with ForceCapture is quasi-static, meaning that at each position, the forces are in static equilibrium, and we only need to compensate for the tool’s gravity. To perform the gravity compensation, we first move ForceCapture in a quasi-static manner for a certain period while recording the pose and wrench data. Using the static equilibrium forces at each position, we construct an overdetermined system of equations to estimate the tool’s center of mass and weight using least-squares solution.
@@ III-C Learning Algorithm: HybridIL
It is important to note that wrench and position control must be orthogonal. While our model does not explicitly model the orthogonality of wrench and position, we achieve this through an orthogonal force-position hybrid controller that aligns with the model’s predicted force-position parameters. This approach differs from conventional imitation learning methods, which typically use a fixed lower-level position controller to track the position commands prediction by the model. HybridIL employs two distinct control primitives to fit the model’s predicted force-position parameters, demonstrated in Fig. 1 (c). When the predicted force is below a threshold of 6N, an IK-based [42] joint position controller is used. If the predicted force exceeds 6N over consecutive steps, a hybrid force-position controller is employed to execute the model’s predicted parameters. The force threshold of 6N was empirically determined. The orthogonal force-position matching approach is illustrated in Fig. 3. For force-position actions where the force exceeds 6N continuously, the motion direction is determined based on the positional information before and after. The corresponding predicted force information is projected onto the orthogonal plane of the motion direction, which defines the force control parameters during execution. For the initial step of hybrid force-position control, if the end-effector has not yet made contact with the object, a pressing control in the opposite direction of force control is applied to achieve stable contact. These functionalities are realized using Flexiv RDK [^1] https://github.com/flexivrobotics/flexiv_rdk of joint position control and hybrid force-position control primitives to execute the force-position actions of HybridIL.
@@ IV Experiments
Figure 4: Experimental setup for data collection efficiency comparison and the time required to fully peel a zucchini by different methods.
@@ Results
The results of the four methods are summarized in TABLE I, and the peeled skins are shown in Fig. 5. The Raw DP method achieved a motion success rate of 80% (16/20), with instances of failure detailed in Fig. 5 (b). Failures marked as ② involved excessive force during the peeling process, which resulted in damage to the zucchini. One instance even broke the bottom of the zucchini, as shown in the bottom of Fig. 5 (b). Failure marked as ④ resulted from no contact with the zucchini, hence no peeling occurred. In contrast, HybridIL demonstrated a 100% success rate (20/20), with all attempts resulting in successful contact and peeling, as illustrated in Fig. 5 (a). When the success criterion was increased to a continuous peeling length of more than 10 cm, both Raw DP and HybridIL experienced a decrease in success rates. Raw DP’s success rate dropped to 55%, with additional failure cases marked as ① and ③. Case ① indicated peeling lengths shorter than 10 cm, and case ③ involved peeling breakage, attributed to discontinuity between the output poses, which caused peeling interruptions. For HybridIL, the success rate decreased to 85%, with failures in cases ① and ③. These failures were due to the premature ending of the output force-position parameters, leading to an early switch from the hybrid force control primitives to IK-based joint position control primitives, which caused peeling discontinuities.
@@
Figure 5: Visualization of the peeled skins by different methods. Failure cases are numbered with circles.
Figure 6: Examples of force curves during peeling a zucchini in different scenarios.
The models that incorporate force as an input, including Force DP and Force+Hybrid DP, performed poorly. While the initial motion leading to the contact with the zucchini was generally correct, once contact occurred, these models struggled to predict the correct pose and force, making successful peeling nearly impossible. This result is counterintuitive—–one would expect that adding force sensing would improve peeling performance, but instead, it worsened the outcome. The reasons for this can be understood from the interactive force curves of the Raw DP and HybridIL during the peeling process, as shown in Fig. 6. Although Raw DP successfully peeled the zucchini, the interactive forces were significantly higher, averaging around 20N and reaching over 40N in some areas. In contrast, the dataset from which these models were trained exhibited much lower interaction forces, around 10N. This mismatch between the input forces and the force distribution in the dataset made it difficult for the models to predict the correct actions. The inconsistency between the force interaction controller used during robot deployment and data collection might be a potential factor contributing to the issue. Addressing this discrepancy could improve the model’s performance. Effectively utilizing the force data collected by ForceCapture as sensory input remains an open challenge and a promising direction for future research. Further exploration on how to better integrate force perception into control strategies could lead to significant advancements in improving task performance. Additionally, as shown in Fig. 5 (a) and Fig. 6 (c), the HybridIL method maintained an average interaction force of 9N, closely aligning with the model’s predicted forces, resulting in evenly peeled sections with consistent thickness and width. However, a slight deviation from the force distribution in the dataset was still noticeable, which likely explains why the Force+Hybrid DP failed to yield good results. Nonetheless, Force+Hybrid DP demonstrated an improvement over the Force DP.
@@ V Conclusion and Discussions
We provide an initial exploration of learning human force-position skills based on ForceCapture. However, there is still room for improvement. 1) Our model uses a simple MLP to represent point clouds, robot pose, and force. In the future, we could explore more advanced multimodal representations that combine visual, force, and robot state data to improve the model’s generalization to diverse skills. 2) HybridIL only employs two control primitives to fit the model’s predicted force-position parameters. Future research could involve exploring more control primitives to better align with model outputs, and the model itself could potentially predict the most suitable primitive and corresponding parameters in advance. 3) ForceMimic has so far demonstrated success with a single peeling skill. In the future, the system could be extended to more force-oriented tasks.
Acknowledgements
This work was supported by the Shanghai Commitee of Science and Technology (No. 24511103200), the National Key Research and Development Project of China (No. 2022ZD0160102), XPLORER PRIZE grants of Shanghai Artificial Intelligence Laboratory, and the National Natural Science Foundation of China (No. 52305030). We would like to thank Flexiv for the hardware of F/T sensor. We are deeply grateful to Shuhan Li, Chen Wang, Wenbo Tang, Jin Liu, Hongjie Fang, Qiaojun Yu, and Qi Wu for their invaluable insights and constructive discussions throughout this research endeavor.